# SAEmnesia: Erasing Concepts in Diffusion Models with Sparse Autoencoders

### v1

@@ 5 Results
Figure 6: Unlearning accuracy before and after UnlearnDiffAtk adversarial attacks. SAEmnesia maintains higher performance under adversarial attack compared to the SAeUron baseline. SAEmnesia drops from 73.30% to 28.30%, compared to SAeUron’s drop from 26.19% to 2.22%.
6 Conclusions and Limitations
We introduce SAEmnesia, a supervised SAE training approach for concept unlearning in diffusion models. Our method mitigates feature splitting and enhances feature centralization, straightening one-to-one concept-latent mappings through automatic concept labeling and direct supervision during training, demonstrating improvements in unlearning effectiveness while maintaining competitive generation quality. Our approach achieves better concept-latent associations compared to unsupervised baselines, reaching 91.51% average performance on the UnlearnCanvas benchmark. Additionally, our methodology reduces hyperparameter search by 96.67% with respect to current approaches, while on sequential unlearning tasks, it demonstrates superior scalability with a 28.4% improvement in unlearning accuracy for 9-object removal. However, we acknowledge that SAEmnesia has some limitations that restrict the generalizability and practical deployment of our approach.
Evaluation Scope. Our evaluation focuses primarily on the UnlearnCanvas benchmark, following the evaluation protocol established in the literature of SAE-based unlearning, to ensure fair comparison. While we employ multiple complementary metrics including unlearning accuracy, retention accuracy across domains, FID scores, and adversarial robustness measures, validation on additional benchmarks with different concept distributions and evaluation protocols would strengthen the generalizability of our findings. Future work could expand evaluation to other concept unlearning benchmarks as they become available in this emerging field.
Computational Trade-offs. Our approach achieves significant computational savings by eliminating the need for percentile-based feature threshold search, which typically requires exploring 1-30+ features per concept in baseline methods. Instead, our supervised training enforces fixed one-neuron-per-concept assignments, substantially reducing the hyperparameter search space. However, our method still requires search over multiplier values, reported in Table 5.
Diffusion Models Architectures. Our experiments focus on SD1.5 [33], following the experimental setup established in the SAE-based unlearning literature, to enable direct performance comparison. While this choice ensures fair evaluation against the state-of-the-art, it limits generalization across the broader landscape of text-to-image generation models. Different diffusion architectures may exhibit varying internal representations, cross-attention mechanisms, and concept encoding strategies that could influence the effectiveness of SAE approaches. Similarly, our focus on specific cross-attention blocks (up.1.1) reflects established practice for concept manipulation in SD1.5, though the optimal intervention points may differ across model architectures. Future work will explore the transferability of our supervised training approach to newer model generations and alternative diffusion architectures.
Ethical Considerations. While our method aims to remove harmful content generation capabilities, concept unlearning techniques could potentially be misused for censorship or may inadvertently encode biases through automatic concept labeling. Due to adversarial attacks, unlearning interventions can be circumvented through carefully crafted prompts, requiring ongoing robustness analysis for safe deployment.

### v2

@@ 4.4 Additional experiments
Figure 9: Unlearning accuracy before and after UnlearnDiffAtk adversarial attacks. SAEmnesia maintains higher performance under adversarial attack compared to the SAeUron baseline.
Nudity unlearning. To highlight the potential of SAEmnesia in real-world applications such as NSFW content removal, we evaluate our method on the I2P benchmark [schramowski2023safe] on SD v1.4. We follow the same experimental setting as cywinski2025saeuron. We train SAEs on SD-v1.4 activations gathered from a random 30K captions from COCO train 2014. We employ the NudeNet detector for nudity detection, filtering out outputs with confidence less than 0.6. Our best model, which steers only the top-2 concept neurons, achieves state-of-the-art results (9 detections vs. SAeUron’s 18), while preserving the model’s overall quality. We adopt two neurons because the SAE is trained on only two nudity-related prompts (“naked man” and “naked woman”). Complete per-category results and comparisons against all baselines are provided in Appendix 6.11.
5 Conclusions
We introduced SAEmnesia, a supervised sparse autoencoder framework for concept unlearning in diffusion models. By preventing feature splitting and enforcing concept-aligned latent structure, SAEmnesia produces more reliable concept–latent mappings and stronger erasure behavior than unsupervised SAE baselines. On UnlearnCanvas, it attains 91.51% average performance while preserving competitive generation quality. The method also reduces hyperparameter search by 96.67%, substantially lowering tuning cost. In sequential unlearning, SAEmnesia scales more effectively, delivering a 28.4% improvement when removing nine objects. We also demonstrate the practical relevance of our approach on nudity removal, where a two-neuron variant of SAEmnesia achieves state-of-the-art NSFW content suppression on I2P while retaining overall image fidelity. Looking forward, as richer benchmarks and regulatory requirements emerge, we see SAEmnesia as a step toward principled, interpretable, and controllable concept unlearning in generative models, with capabilities increasingly critical for trustworthy and safe deployment.
Acknowledgments We acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support.
6 Appendix
6.1 Standard deviations
We report here the standard deviations of UA, IRA and CRA metrics across 4 different seeds.
Table 2: Mean and standard deviations (%) of SAEmnesia on the UnlearnCanvas benchmark.
Method UA IRA CRA Avg.
SAEmnesia 94.65 $\pm$ 2.6 91.39 $\pm$ 1.3 88.48 $\pm$ 0.5 91.51 $\pm$ 0.4
6.2 Multipliers Comparison
The comparison reveals significant differences in unlearning strategies between other SAE-based methods [cywinski2025saeuron] and our SAEmnesia variants. SAeUron employs substantially larger multiplier magnitudes (as per Tab. 3), averaging -21.25, while our SAEmnesia-OS-CA-FS and SAEmnesia-OS-DC-CA-FT models achieve effective unlearning with more conservative multipliers, averaging -6.20 and -6.60 respectively. These results align well with the findings on latents score distributions shown in Figs. 4, 10, 11 and 12.
Table 3: Comparison of Object Unlearning Multipliers Across Three Models
Object SAeUron S-OS-CA-FS S-OS-DC-CA-FT
Architectures -20.0 -5.0 -5.0
Bears -30.0 -10.0 -5.0
Birds -10.0 -5.0 -5.0
Butterfly -15.0 -5.0 -5.0
Cats -15.0 -1.0 -10.0
Dogs -20.0 -5.0 -5.0
Fishes -30.0 -5.0 -5.0
Flame -25.0 -1.0 -1.0
Flowers -20.0 -5.0 -5.0
Frogs -5.0 -5.0 -10.0
Horses -25.0 -10.0 -15.0
Human -20.0 -5.0 -5.0
Jellyfish -15.0 -1.0 -1.0
Rabbits -30.0 -10.0 -5.0
Sandwiches -15.0 -25.0 -5.0
Sea -30.0 -5.0 -5.0
Statues -30.0 -1.0 -10.0
Towers -20.0 -5.0 -5.0
Trees -25.0 -5.0 -5.0
Waterfalls -30.0 -10.0 -20.0
Average -21.25 $\pm$ 7.45 -6.20 $\pm$ 5.31 -6.60 $\pm$ 4.48
6.3 Additional Scores Histograms
In this section are provided additional examples similar to Fig. 4, to further demonstrate the robustness of our methods throughout different concepts, such as Architectures Fig. 10, Rabbits Fig. 11 and Sea Fig. 12. For all the objects, we show a clear improvement, going from distributed scores (in orange) produced by SAeUron [cywinski2025saeuron], to well defined peaks (in green) produced by our method SAEmnesia.
Figure 10: Feature importance score distributions for Architectures concept. SAeUron shows dispersed concept scores, while SAEmnesia shows a clear dominant peak.
Figure 11: Feature importance score distributions for Rabbits concept. SAeUron shows dispersed concept scores, while SAEmnesia shows a clear dominant peak.
Figure 12: Feature importance score distributions for Sea concept. SAeUron shows dispersed concept scores, while SAEmnesia shows a clear dominant peak.
6.4 Additional unlearning visualization
Figs. 13 and 14 provides additional visualization with randomly sampled styles when unlearning with SAEmnesia. We can see that SAEmnesia performs well in most cases; however, it struggles in some scenarios with more challenging objects that are more blended with the styles, for example, flames and jellyfishes.
Figs. 15 and 16 show the unlearning accuracies for each object in the columns when unlearning the objects in the rows. This visualization highlights entanglements between different objects that can lead to poor unlearning performances.
Figure 13: Qualitative examples of concept removal with SAEmnesia. Each row shows a different randomly sampled style, and each column a different object concept.
Figure 14: Qualitative examples of concept removal with SAEmnesia. Each row shows a different randomly sampled style, and each column a different object concept.
6.5 Global Cross-Entropy
We employ Cross-entropy (CE) loss with softmax applied to all pre-TopK latent activations as an alternative to the Concept-Assignment Loss (Eq. 8). This way, we treat concept assignment as a classification problem across the entire latent space:
$\mathcal{L}_{\text{global-ce}}=-\frac{1}{B}\sum_{i=1}^{B}\log\left(\frac{\exp(v^{i}_{c_{i}})}{\sum_{j=1}^{n}\exp(v^{i}_{j})}\right),$ (11)
where $B$ is the batch size, $v^{i}_{j}$ represents the pre-TopK activation value for sample $i$ at latent position $j$, $c_{i}$ denotes the assigned concept latent index for sample $i$, and $n$ is the total number of latents.
This approach, however, led to suboptimal performance as reported in Tabs. 5 and 6 w.r.t. $O$-GCE-FS and $O$-GCE-FT.
6.6 Post-TopK Loss Analysis
OS-TK-CA-FT is the only model variant that applies CA loss after the Top-K, yet demonstrates competitive performance without major degradation. In fine-tuned configurations with hyperparameter search (Tab. 5), OS-TK-CA-FT achieves 85.52% average performance compared to the baseline’s 82.29%, representing a meaningful 3.2% improvement. Similarly, in from-scratch training (Tab. 6), OS-TK-CA-FT maintains 86.93% performance, indicating that post Top-K supervision remains viable.
6.7 Exploratory Variants
In Tab. 4, all the variants tested for this work are presented. SAEmnesia-OS-DC-CA-FT is the version reported as SAEmnesia in the main paper.
Table 4: SAE Experiment Versions.
Feature/Metric $O$-GCE-FS $O$-GCE-FT OS-DC-CA-FS OS-DC-CA-FT $O$-CA-FS $O$-CA-FT OS-CA-FS OS-CA-FT OS-TK-CA-FS OS-TK-CA-FT
Object labels
Styles labels $\times$ $\times$ $\times$ $\times$
Decorrelation $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Global CE $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
CA before Top-K $\times$ $\times$ $\times$ $\times$
CA after Top-K $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Finetuned $\times$ $\times$ $\times$ $\times$ $\times$
Table 5: SAEmnesia Experimental Results - Fine Tuned Models - Searched HP
Metric SAeUron $O$-GCE-FT OS-DC-CA-FT $O$-CA-FT OS-CA-FT OS-TK-CA-FT
UA (%) $\uparrow$ 87.16 94.55 91.75 95.75 90.90 66.85
IRA (%) $\uparrow$ 85.57 87.44 93.16 85.92 92.11 96.76
CRA (%) $\uparrow$ 74.14 57.71 88.60 75.09 86.46 92.95
Avg. (%) $\uparrow$ 82.29 79.9 91.51 85.59 89.82 85.52
FID (%) $\downarrow$ 124.11 155.17 111.16 111.70 110.84 110.24
Table 6: SAEmnesia Experimental Results - Trained From Scratch with Hyperparameters Search
Metric SAeUron $O$-GCE-FS OS-DC-CA-FS $O$-CA-FS OS-CA-FS OS-TK-CA-FS
UA (%) $\uparrow$ 87.16 95.98 92.95 78.45 86.00 69.45
IRA (%) $\uparrow$ 85.57 68.20 91.48 64.25 92.76 97.38
CRA (%) $\uparrow$ 74.14 42.48 88.12 57.76 87.05 93.97
Avg. (%) $\uparrow$ 82.29 69.22 90.85 66.82 88.93 86.93
FID (%) $\downarrow$ 124.11 150.15 110.25 119.46 110.49 109.84
6.8 Concept Interference Mitigation
The Dogs vs. Cats overlap analysis (Tabs. 7 and 8) provides crucial insights into concept interference patterns, a known limitation of the current approaches [cywinski2025saeuron].
From-Scratch Models generally show better concept separation, with OO-GCE-FS achieving zero overlapping timesteps.
Fine-Tuned Models show more variable performance, with some variants (OO-GCE-FT, OS-TK-CA-FT) achieving zero overlap while others (OS-CA-FT) perform worse than the baseline.
Table 7: Timesteps with Cats-Dogs Overlapping Latent as Most Active - Models Trained from scratch
Metric SAeUron $O$-GCE-FS OS-DC-CA-FS $O$-CA-FS OS-CA-FS OS-TK-CA-FS
Overlap. timesteps 10 0 3 6 10 2
Table 8: Timesteps with Cats-Dogs Overlapping Latent as Most Active - Finetuned Models
Metric SAeUron $O$-GCE-FT OS-DC-CA-FT $O$-CA-FT OS-CA-FT OS-TK-CA-FT
Overlap. timesteps 10 0 10 3 14 0
Figure 15: OS-DC-CA-FT UA and IRA classes disentanglement.
Figure 16: OS-CA-FS UA and IRA classes disentanglement.
6.9 Uniform Multipliers Analysis
The uniform multiplier sweep analysis (Figs. 17 and 18) reveals varying performance characteristics across model versions and evaluation metrics. In the fine-tuned models (Fig. 17), SAEmnesia variants show competitive performance with some variations across different multiplier ranges, with certain variants like S-OS-DC-CA-FT maintaining strong performance while others such as S-OS-TK-CA-FT exhibit sensitivity to specific multiplier settings. The from-scratch models (Fig. 18) demonstrate different behavior patterns, with variants like S-OO-CA-FS showing particular sensitivity to moderate multiplier values before recovering at gentler settings. Both figures indicate that SAEmnesia variants achieve reasonable performance across various hyperparameter ranges. The results suggest that different SAEmnesia configurations may be better suited for different multiplier ranges, highlighting the importance of hyperparameter selection in optimizing unlearning performance.
Figure 17: All Model’s performances with uniform multipliers - fine tuned.
Figure 18: All Model’s performances with uniform multipliers - from scratch.
6.10 Lambda Coefficient Analysis for SAEmnesia Loss
The choice of value for $\beta$ in Eq. 6 as the coefficient for the $\mathcal{L}_{supSAE}$ loss critically affects model performance. Our best experimental results were achieved with $\beta=3$. When $\beta=10$, the model produces completely white images and loses generative capabilities. In our experimental setup, $\lambda$ from Eq. 6 was set to $\lambda=0.01$, and $\gamma$ from Eq. 9 was set to $\gamma=0.1$.
6.11 Nudity Unlearning
Method Armpits Belly Buttocks Feet Breasts (F) Genitalia (F) Breasts (M) Genitalia (M) Total CLIPScore (↑) FID (↓)
FMN [zhang2024forget] 43 117 12 59 155 17 19 2 424 30.39 13.52
CA [kumari2023ablating] 153 180 45 66 298 22 67 7 838 31.37 16.25
AdvUn [zhang2024defensive] 8 0 0 13 1 1 0 0 28 28.14 17.18
Receler [huang2024receler] 48 32 3 35 20 0 17 5 160 30.49 15.32
MACE [lu2024mace] 17 19 2 39 16 0 9 7 111 29.41 13.42
CPE [lee2025concept] 10 8 2 8 6 1 3 2 40 31.19 13.89
UCE [gandikota2024unified] 29 62 7 29 35 5 11 4 182 30.85 14.07
SLD-M [schramowski2023safe] 47 72 3 21 39 1 26 3 212 30.90 16.34
ESD-x [gandikota2023erasing] 59 73 12 39 100 6 18 8 315 30.69 14.41
ESD-u [gandikota2023erasing] 32 30 2 19 27 3 8 2 123 30.21 15.10
SAeUron [cywinski2025saeuron] 7 1 3 2 4 0 0 1 18 30.89 14.37
SAEmnesia 7 17 2 5 11 2 2 1 47 30.98 14.72
SAEmnesia-top2 1 3 1 0 4 0 0 0 9 30.98 14.72
SD v1.4 148 170 29 63 266 18 42 7 743 31.34 14.04
SD v2.1 105 159 17 60 177 9 57 2 586 31.53 14.87
Table 9: SAEmnesia-top2 is the presented SAEmnesia setup, but with two latents steered instead of one. Multiplier for SAEmnesia (one latent affected): -49. Multiplier for SAEmnesia-top2 (two latents affected): -60.
For nudity unlearning, we follow the same experimental setting as cywinski2025saeuron as described in the main paper. Our base unlearning setup uses only the top scoring latent to erase an object. As per Tab. 9, for nudity unlearning, SAEmnesia achieves weak performance (47 detections vs. SAeUron’s 18 detections). However, when we instead steer two latents in our variant SAEmnesia-top2, performance substantially improves (9 detections). The highly imbalanced distribution of nudity-related content in the training dataset may lead to a weaker concept centralization.
Nevertheless, SAEmnesia maintains a significant practical advantage over unsupervised alternatives. SAeUron requires steering the 205 top-scoring latents to achieve such performances, while SAEmnesia-top2 only needs 2.